# The vertical structure of debris discs and the role of disc gravity

### v1

@@ Abstract
Debris discs provide valuable insights into the formation and evolution of exoplanetary systems. Their structures are commonly attributed to planetary perturbations, serving as probes of as-yet-undetected planets. However, most studies of planet–debris disc interactions ignore the disc’s gravity, treating it as a collection of massless planetesimals. Here, using an analytical model, we investigate how the vertical structure of a back-reacting debris disc responds to secular perturbations from an inner, inclined planet. Considering the disc’s axisymmetric potential, we identify two dynamical regimes: planet-dominated and disc-dominated, which may coexist, separated by a secular-inclination resonance. In the planet-dominated regime ($M_{d}/m_{p}\ll 1$), we recover the classical result: a transient warp propagates outward until the disc settles into a box-like structure centered around the planetary orbit’s initial inclination $I_{p}(0)$, with a distance-independent aspect ratio $\mathcal{H}(R)\approx I_{p}(0)$. In contrast, in the disc-dominated regime ($M_{d}/m_{p}\gtrsim 1$), the disc exhibits dynamical rigidity, remaining thin and misaligned, with significantly suppressed inclinations and a sharply declining aspect ratio, $\mathcal{H}(R)\propto I_{p}(0)R^{-7/2}$. In the intermediate regime ($M_{d}/m_{p}\lesssim 1$), the system exhibits a secular-inclination resonance, leading to long-lived, warp-like structures and a bimodal inclination distribution, containing both dynamically hot and cold populations. We provide analytic formulae describing these effects as a function of planet-disc parameters. We also find that the vertical density profile is intrinsically non-Gaussian and recommend fitting disc observations with non-zero slopes of $\mathcal{H}(R)$. Our results may be used to infer planetary parameters and debris disc masses based on observed warps and scale heights, as demonstrated for HD 110058 and $\beta$ Pic.
@@
pubyear: 2025pagerange: The vertical structure of debris discs and the role of disc gravityD.3
@@ 1.1 Motivation
To date, around 20 per cent of main sequence stars in the solar neighborhood are known to host debris disks (Montesinos et al., 2016; Sibthorpe et al., 2018). These circumstellar disks, representing remnants of planet-formation processes, are typically optically thin and gas-poor (Hughes et al., 2018; Wyatt, 2020), although a few contain detectable amounts of gas (e.g., Marino et al., 2016; Kral et al., 2020). Debris disks are populated by solids ranging from micron-sized dust grains to kilometer-sized or larger planetesimals, akin to the asteroid and Kuiper belts in the Solar System. Unlike the primordial interstellar dust in protoplanetary disks, debris disk dust is considered to be secondary in origin, continuously replenished through collisional cascades among larger parent planetesimals (Wyatt, 2008). Thus, sustaining the levels of dust estimated from observations – typically, $\sim 10^{-3}-1M_{\earth}$ (Wyatt & Dent, 2002; Panić et al., 2013; Holland et al., 2017; Matrà et al., 2025) – requires a massive reservoir of parent planetesimals. Depending on the assumed properties, such as size distribution and collisional behaviors, estimates for this reservoir can be as high as $\sim 10^{3}M_{\earth}$ (Krivov & Wyatt, 2021). As such, debris disks are often considered massive analogues of the Kuiper belt (Wyatt, 2020).
@@
Vertical structures in particular, including warps and scale heights (or aspect ratios), are of central interest. Indeed, warped structures – characterized by a tilted inner disk extending into a flat outer disk – are often interpreted as evidence of planets on inclined orbits (e.g., Augereau et al., 2001; Dawson et al., 2011; Nesvold & Kuchner, 2015; Nesvold et al., 2017; Brady et al., 2023; Farhat et al., 2023; Smallwood, 2023). Part of the argument here is that debris disks are thought to settle into nearly flat configurations once the protoplanetary disk dissipates, in the absence of other perturbers. An iconic example is the $\beta$ Pic system, where the detection of a warped structure (Lagage & Pantin, 1994; Kalas & Jewitt, 1995) led to the prediction (Mouillet et al., 1997; Augereau et al., 2001) and subsequent direct imaging of the planet responsible for it (Lagrange et al., 2010), $\beta$ Pic b. [^1] It is noteworthy that $\beta$ Pic b is not completely aligned with the inner disk as expected based on theoretical considerations; for a detailed discussion, see Dawson et al. (2011). Additionally, a second inner planet, $\beta$ Pic c, has been detected (Lagrange et al., 2019); this, however, does not appear to significantly affect the warp (Dong et al., 2020; Smallwood, 2023). While warps are most readily accessible in disks with near edge-on configurations – introducing observational biases against their detection – a few systems have been identified to date: namely, AU Mic (Wang et al., 2015), HD 111520 (Draper et al., 2016; Crotts et al., 2022), and HD 110058 (Kasper et al., 2015; Hales et al., 2022; Stasevic et al., 2023).
Another observationally measurable vertical property of debris disks is their scale height and thus aspect ratio $\mathcal{H}$. Such measurements have been made for tens of debris disks, revealing a striking diversity, with values ranging from $\mathcal{H}\sim 0.01$ to $0.3$ (e.g., Kennedy et al., 2018; Hales et al., 2022; Marshall et al., 2023; Terrill et al., 2023; Matrà et al., 2025). Interestingly, while most measurements assume a radius-independent aspect ratio (to simplify fitting procedures), advanced (non-parametric) fitting methods have also revealed radius-dependent variations (Han et al., 2022, 2025). Additionally, wavelength-dependent trends are beginning to emerge, adding further complexity. Indeed, while most measured aspect ratios are consistent across wavelengths, observations of the AU Mic debris disk suggest an increasing aspect ratio with wavelength (Daley et al., 2019; Olofsson et al., 2022; Vizgan et al., 2022). Finally, a recent study analyzing debris disks observed with ALMA suggests a tentative trend of a decreasing aspect ratio with increasing dust mass around A-type stars (Han et al., 2025).
@@ 1.2 Previous works
Three key mechanisms have been explored in the literature to explain vertical structures in debris disks. The first mechanism involves flybys, where the passage of a nearby star or massive object can induce tilts in the disk’s vertical structure, as well as puff it up (Lestrade et al., 2011; Moore et al., 2020; Batygin et al., 2020). The second mechanism involves gravitational stirring by large bodies embedded within the disk itself–such as Pluto-sized or larger embryos–where two-body encounters with these bodies dynamically excite the surrounding smaller planetesimals (e.g., Ida, 1990; Ida & Makino, 1992, 1993; Kokubo & Ida, 2012). This mechanism can thicken the disk vertically (e.g., Muñoz-Gutiérrez et al., 2015; Daley et al., 2019; Matrà et al., 2019), but when operating alone, it is not expected to induce warps. Recently, however, the combined effects of this mechanism and a planet located interior to the disk – referred to as ‘mixed-stirring’ – have also been explored (Muñoz-Gutiérrez et al., 2023).
@@ 2 Model System
[…]
The debris disk, on the other hand, extending from $a_{\rm in}$ to $a_{\rm out}$, consists of $N\gg 1$ planetesimals on circular orbits. Each planetesimal is characterized by its mass $m_{j}$, orbital semimajor axis $a_{d}=\{a_{j}\}$, inclination $I_{d}=\{I_{j}\}$, longitude of ascending node $\Omega_{d}=\{\Omega_{j}\}$, and mean motion $n_{d}=\{n_{j}\}\approx\sqrt{GM_{c}/a_{j}^{3}}$ (with $j=1,\ldots,N$). The disk is taken to be initially razor-thin with $I_{j}(0)=0$ and $\Omega_{j}(0)$ that is uniformly drawn between 00 and $2\pi$. The semimajor axes of the planetesimals are taken to be logarithmically spaced, with a geometric spacing ratio of $a_{j+1}/a_{j}=(a_{\rm out}/a_{\rm in})^{1/{(N-1)}}$. Based on this, the masses $m_{j}$ are assigned such that the disk’s initial surface-density distribution follows a truncated power-law profile given by:
@@ 3.1.1 Effects of the Planet on Planetesimals
[…]
@@ 3.1.2 Effects of Massive Planetesimals on the Planet
where the numerical estimate is obtained for the fiducial disk model ($p=3/2$, $\delta=5$), assuming $a_{p}/a_{\rm in}=2/3$ so that $\phi_{1}^{c}\approx 1.85$ (more on this below). A similar expression can also be derived for the summation involving the $\delta B_{j,p}$ term in Equation (8), provided that certain assumptions are made about $I_{j}$ and $\Delta\Omega\equiv\Omega_{j}-\Omega_{p}$, which may or may not depend on $a_{j}$. Such a calculation is presented in Appendix A; however, as this paper neglects the disk’s non-axisymmetric torque, we do not elaborate further on its behavior.
@@
Figure 2: The ratio $A_{d,p}/A_{p}$, i.e., the disk-induced planetary nodal recession rate ($A_{d,p}<0$; Eq. (9)) relative to that of the planetesimals due to the planet ($A_{p}(a_{d})<0$; Eq. (3.1.1)), as a function of planetesimal semimajor axis $a_{d}/a_{\rm in}$. Calculations are done for various disk-to-planet mass ratios, $M_{d}/m_{p}$ (shown by different colors), assuming the fiducial disk model ($p=3/2$, $a_{\rm in}=30$ au, $\delta=5$; Section 2) and a planet at $a_{p}=20$ au. For $10^{-2}\lesssim M_{d}/m_{p}\lesssim 10$, the nodal recession rates are equal, $A_{d,p}/A_{p}=1$, at a unique semimajor axis $a_{\rm res}$. Interior to this, $|A_{d,p}|\lesssim|A_{p}|$, while exterior to it, $|A_{d,p}|\gtrsim|A_{p}|$. For more massive disks, i.e., $M_{d}/m_{p}\gtrsim 10$, the ratio $A_{d,p}/A_{p}\gtrsim 1$ across the entire disk. Conversely, for low-mass disks, i.e., $M_{d}/m_{p}\lesssim 10^{-2}$, the ratio $A_{d,p}/A_{p}\lesssim 1$ at all semimajor axes. Note that $A_{d,p}/A_{p}\propto a_{d}^{7/2}$ (black dotted line; Eq. 4.1). See the text (Sections 3.1 and 4.1) for more details.
@@ 3.2 General Evolution Equations
where Equation (13) applies for $j=1,...,N$. These $N+1$ coupled ordinary differential equations can be further compacted into a matrix equation of the form $\dot{\zeta}=i\mathbf{B}\zeta$:
@@
where $a_{1}=a_{\rm in}$ and $a_{N}=a_{\rm out}$. Equation (15) represents the master equation needed for our work, encapsulating the coupled secular inclination evolution of the planet and the back-reacting disk. In this work, we neglect the terms $\delta B_{j,p}$ in Equation (15); consideration of this non-axisymmetric component is deferred to future work (see Section 6.4). As we shall see, this allows us to fully flesh out the basic effects of disk gravity analytically. Finally, we note that in the case of a massless disk, $A_{d,p}=\delta B_{j,p}=0$ for all $j=1,..,N$, and Equation (15) reduces to the classical evolution equations of test-particles due to an internal, non-precessing planet on an inclined orbit (Murray & Dermott, 1999). Similarly, for $N=1$, one recovers the equations governing the secular interactions of two mutually inclined planets on circular orbits.
4 Secular behavior: analytical results
We aim to elucidate the dynamical effects of the disk’s back-reaction on the planet, providing insights into the vertical structure of massive debris disks. Although Equation (15) does not admit an analytical closed-form solution for $M_{d}\neq 0$, [^5] The difficulty arises due to the non-axisymmetric component of the disk’s potential which is proportional to $I_{p}I_{d}\cos\Delta\Omega$ and may vary over time (Equation 8). In principle, an approximate analytical solution could be attempted with an educated guess for the time-averaged $I_{d}(t)$ and $\Delta\Omega(t)$, but this is not obvious to us a priori (see also Appendix B). progress can be made in a simplified case by considering only the axisymmetric component of the disk’s potential (i.e., setting $\delta B_{j,p}=0$ for all $j$), as outlined below.
In this simplified case, assuming planetesimals are initiated on coplanar orbits ($\zeta_{j}(0)=0$), Equation (15) can be solved analytically. Doing so, we find that the planetary orbit evolves such that:

### v2

@@ Abstract
Debris discs provide valuable insights into the formation and evolution of exoplanetary systems. Their structures are commonly attributed to planetary perturbations, serving as probes of as-yet-undetected planets. However, most studies of planet-debris disc interactions ignore the disc’s gravity, treating it as a collection of massless planetesimals. We develop a simplified analytical model as a primer to investigate how the vertical structure of a back-reacting debris disc responds to secular perturbations from an inner, inclined planet. Considering the disc’s axisymmetric potential, we identify two dynamical regimes: planet-dominated and disc-dominated, which may coexist, separated by a secular-inclination resonance. In the planet-dominated regime ($M_{d}/m_{p}\ll 1$), we recover the classical result: a transient warp propagates outward until the disc settles into a box-like structure centered around the planetary orbit’s initial inclination $I_{p}(0)$, with a distance-independent aspect ratio $\mathcal{H}(R)\approx I_{p}(0)$. In contrast, in the disc-dominated regime ($M_{d}/m_{p}\gtrsim 1$), the disc exhibits dynamical rigidity, remaining thin and misaligned, with significantly suppressed inclinations and a sharply declining aspect ratio, $\mathcal{H}(R)\propto I_{p}(0)R^{-7/2}$. In the intermediate regime ($M_{d}/m_{p}\lesssim 1$), the system exhibits a secular-inclination resonance, leading to long-lived, warp-like structures and a bimodal inclination distribution, containing both dynamically hot and cold populations. We provide analytic formulae describing these effects as a function of system parameters. We also find that the vertical density profile is intrinsically non-Gaussian and recommend fitting observations with non-zero slopes of $\mathcal{H}(R)$. Our results may be used to infer planetary parameters and debris disc masses based on observed warps and scale heights, as demonstrated for HD 110058 and $\beta$ Pic.
@@
pubyear: 2025pagerange: The vertical structure of debris discs and the role of disc gravity: A primer using a simplified modelD.3
@@ 1.1 Motivation
To date, around 20 per cent of main sequence stars in the solar neighborhood are known to host debris disks (Montesinos et al., 2016; Sibthorpe et al., 2018). These circumstellar disks, representing remnants of planet-formation processes, are typically optically thin and gas-poor (Hughes et al., 2018; Wyatt, 2020), although a few contain detectable amounts of gas (e.g., Marino et al., 2016; Kral et al., 2020). Debris disks are populated by solids ranging from observed micron-sized dust grains to inferred kilometer-sized or larger planetesimals, akin to the asteroid and Kuiper belts in the Solar System. Unlike the primordial interstellar dust in protoplanetary disks, debris disk dust is considered to be secondary in origin, continuously replenished through collisional cascades among larger parent planetesimals (Wyatt, 2008). Thus, sustaining the levels of dust estimated from observations – typically, $\sim 10^{-3}-1M_{\earth}$ (Wyatt & Dent, 2002; Panić et al., 2013; Holland et al., 2017; Matrà et al., 2025) – requires a massive reservoir of parent planetesimals. Depending on the assumed properties, such as size distribution and collisional behaviors, estimates for this reservoir can be as high as $\sim 10^{3}M_{\earth}$ (Krivov & Wyatt, 2021). As such, debris disks are often considered massive analogues of the Kuiper belt (Wyatt, 2020).
@@
Vertical structures in particular, including warps and scale heights (or aspect ratios), are of central interest. Indeed, warped structures – characterized by a tilted inner disk extending into a flat outer disk – are often interpreted as evidence of planets on inclined orbits (e.g., Augereau et al., 2001; Dawson et al., 2011; Nesvold & Kuchner, 2015; Nesvold et al., 2017; Brady et al., 2023; Farhat et al., 2023; Smallwood, 2023). Part of the argument here is that debris disks are expected to retain the nearly flat configurations established during the gas-rich protoplanetary disk phase, in the absence of other perturbers. An iconic example is the $\beta$ Pic system, where the detection of a warped structure (Lagage & Pantin, 1994; Kalas & Jewitt, 1995) led to the prediction (Mouillet et al., 1997; Augereau et al., 2001) and subsequent direct imaging of the planet responsible for it (Lagrange et al., 2010), $\beta$ Pic b. [^1] It is noteworthy that $\beta$ Pic b is not completely aligned with the inner disk as expected based on theoretical considerations; for a detailed discussion, see Dawson et al. (2011). Additionally, a second inner planet, $\beta$ Pic c, has been detected (Lagrange et al., 2019); this, however, does not appear to significantly affect the warp (Dong et al., 2020; Smallwood, 2023). While warps are most readily accessible in disks with near edge-on configurations – introducing observational biases against their detection – a few systems have been identified to date: namely, AU Mic (Wang et al., 2015), HD 111520 (Draper et al., 2016; Crotts et al., 2022), and HD 110058 (Kasper et al., 2015; Hales et al., 2022; Stasevic et al., 2023).
Another observationally measurable vertical property of debris disks is their scale height $\mathfrak{h}_{d}$ and thus aspect ratio $\mathcal{H}$ $\equiv\mathfrak{h}_{d}/R$. Such measurements have been made for tens of debris disks, revealing a striking diversity, with aspect ratio values ranging from $\sim 0.01$ to $0.3$ (e.g., Kennedy et al., 2018; Hales et al., 2022; Marshall et al., 2023; Terrill et al., 2023; Matrà et al., 2025). Interestingly, while most measurements assume a radius-independent aspect ratio (to simplify fitting procedures), advanced (non-parametric) fitting methods have also revealed radius-dependent variations (Han et al., 2022, 2025). Additionally, observations of the AU Mic debris disk suggest a wavelength-dependent increase in aspect ratio (Daley et al., 2019; Olofsson et al., 2022; Vizgan et al., 2022). While AU Mic is currently the only system with such evidence, it highlights the potential for wavelength-dependent aspect ratios, adding further complexity. Finally, a recent study analyzing debris disks observed with ALMA suggests a tentative trend of a decreasing aspect ratio with increasing dust mass around A-type stars (Han et al., 2025).
@@ 1.2 Previous works
Three key mechanisms have been explored in the literature to explain vertical structures in debris disks. The first mechanism involves flybys, where the passage of a nearby star or massive object can induce tilts in the disk’s vertical structure, as well as puff it up (Kobayashi & Ida, 2001; Lestrade et al., 2011; Moore et al., 2020; Batygin et al., 2020). The second mechanism involves gravitational stirring by large bodies embedded within the disk itself–such as Pluto-sized or larger embryos–where two-body encounters with these bodies dynamically excite the surrounding smaller planetesimals (e.g., Ida, 1990; Ida & Makino, 1992, 1993; Kokubo & Ida, 2012). This mechanism can thicken the disk vertically (e.g., Muñoz-Gutiérrez et al., 2015; Daley et al., 2019; Matrà et al., 2019), but when operating alone, it is not expected to induce warps. Recently, however, the combined effects of this mechanism and a planet located interior to the disk – referred to as ‘mixed-stirring’ – have also been explored (Muñoz-Gutiérrez et al., 2023).
@@ 2 Model System
We characterize the planetary orbit, assumed to be circular, by its semimajor axis $a_{p}$, inclination $I_{p}$, longitude of ascending node $\Omega_{p}$, and mean motion $n_{p}\approx\sqrt{GM_{c}/a_{p}^{3}}$. The planet is taken to be initially inclined with respect to the disk (which lies in the equatorial plane of the star, more on this below), typically with $I_{p}(0)\lesssim 20^{\circ}$ and, without loss of generality, $\Omega_{p}(0)=0$. [^2] We limit $I_{p}(0)$ for two reasons: first, the linear Laplace-Lagrange secular theory we employ (Section 3) is valid for small inclinations; second, to avoid scenarios where von Zeipel-Lidov-Kozai effects (Lidov, 1962; Kozai, 1962; Naoz, 2016; Ito & Ohtsuka, 2019) cannot be ignored. The latter can modulate the eccentricities of the bodies involved, introducing further complexity.
The debris disk, on the other hand, extending from $a_{\rm in}$ to $a_{\rm out}$, consists of $N\gg 1$ planetesimals on circular orbits. Each planetesimal is characterized by its mass $m_{j}$, orbital semimajor axis $a_{d}=\{a_{j}\}$, inclination $I_{d}=\{I_{j}\}$, longitude of ascending node $\Omega_{d}=\{\Omega_{j}\}$, and mean motion $n_{d}=\{n_{j}\}\approx\sqrt{GM_{c}/a_{j}^{3}}$ (with $j=1,\ldots,N$). The disk is taken to be initially razor-thin with $I_{j}(0)=0$ and $\Omega_{j}(0)$ that is uniformly drawn between 0 and $2\pi$. The semimajor axes of the planetesimals are taken to be logarithmically spaced, with a geometric spacing ratio of $a_{j+1}/a_{j}=(a_{\rm out}/a_{\rm in})^{1/{(N-1)}}$. Based on this, the masses $m_{j}$ are assigned such that the disk’s initial surface-density distribution follows a truncated power-law profile given by:
@@
Here, we note in passing that Laplace-Lagrange theory for secular inclination dynamics (as adopted in this work; Section 3) does not prescribe a natural reference plane. As a result, it is most meaningful to consider mutual inclinations, rendering the choice of a reference plane arbitrary. In this work, we adopt the stellar equatorial plane as our reference (Fig. 1).
@@ 3.1.1 Effects of the Planet on Planetesimals
$b_{s}^{(m)}(\alpha)=\frac{2}{\pi}\int_{0}^{\pi}\cos(m\theta)\big{[}1+\alpha^{2}-2\alpha\cos\theta\big{]}^{-s}d\theta,$ (7)
@@ 3.1.2 Effects of Massive Planetesimals on the Planet
where the numerical estimate is obtained for the fiducial disk model ($p=3/2$, $\delta=5$), assuming $a_{p}/a_{\rm in}=2/3$ so that $\phi_{1}^{c}\approx 1.85$. Here, $\phi_{1}^{c}$ is a dimensionless factor of order unity accounting for the contributions of planetesimals close to the planetary orbit (more on this below and in App. A). A similar expression can also be derived for the summation involving the $\delta B_{j,p}$ term in Equation (8), provided that certain assumptions are made about $I_{j}$ and $\Delta\Omega\equiv\Omega_{j}-\Omega_{p}$, which may or may not depend on $a_{j}$. Such a calculation is presented in Appendix A; however, as this paper neglects the disk’s non-axisymmetric torque, we do not elaborate further on its behavior.
@@
Figure 2: The ratio $A_{d,p}/A_{p}$, i.e., the disk-induced planetary nodal recession rate ($A_{d,p}<0$; Eq. (9)) relative to that of the planetesimals due to the planet ($A_{p}(a_{d})<0$; Eq. (3.1.1)), as a function of planetesimal semimajor axis $a_{d}/a_{\rm in}$. Calculations are done for various disk-to-planet mass ratios, $M_{d}/m_{p}$ (shown by different colors), assuming the fiducial disk model ($p=3/2$, $a_{\rm in}=30$ au, $\delta=5$; Section 2) and a planet at $a_{p}=20$ au. For $10^{-2}\lesssim M_{d}/m_{p}\lesssim 10$, the nodal recession rates are equal, $A_{d,p}/A_{p}=1$, at a unique semimajor axis $a_{\rm res}$. Interior to this, $|A_{d,p}|\lesssim|A_{p}|$, while exterior to it, $|A_{d,p}|\gtrsim|A_{p}|$. For more massive disks, i.e., $M_{d}/m_{p}\gtrsim 10$, the ratio $A_{d,p}/A_{p}\gtrsim 1$ across the entire disk. Conversely, for low-mass disks, i.e., $M_{d}/m_{p}\lesssim 10^{-2}$, the ratio $A_{d,p}/A_{p}\lesssim 1$ at all semimajor axes. Note that $A_{d,p}/A_{p}\propto a_{d}^{7/2}$ for $a_{d}/a_{\rm in}\gg 1$ (black dotted line; Eq. 4.1). See the text (Sections 3.1 and 4.1) for more details.
@@ 3.2 General Evolution Equations
where Equation (13) applies for j=1,,Nj=1,...,N. These $N+1$ coupled ordinary differential equations can be further compacted into a matrix equation of the form $\dot{\zeta}=i\mathbf{B}\zeta$:
@@
where $a_{1}=a_{\rm in}$ and $a_{N}=a_{\rm out}$. Equation (15) represents the master equation needed for our work, encapsulating the coupled secular inclination evolution of the planet and the back-reacting disk. In this work, we neglect the terms $\delta B_{j,p}$ in Equation (15); full consideration of this non-axisymmetric component is deferred to future work (see, however, Section 6.4 and App. B). As we shall see, this allows us to fully flesh out the basic effects of disk gravity analytically.
Finally, we note that in the case of a massless disk, $A_{d,p}=\delta B_{j,p}=0$ for all $j=1,..,N$, and Equation (15) reduces to the classical evolution equations of test-particles due to an internal, non-precessing planet on an inclined orbit (Murray & Dermott, 1999). Similarly, for $N=1$, one recovers the equations governing the secular interactions of two mutually inclined planets on circular orbits. Note that this is because when $N=1$, the planetesimal is formally treated as a single non-self-gravitating ring. To model a self-gravitating ring instead, one must consider $N>1$ planetesimals sharing a fixed, single-valued semimajor axis and spatially soften the ring-ring interaction potential – e.g., by modeling the planetesimals as Plummer spheres – to properly account for their mutual interactions (Touma, 2002; Hahn, 2003; Sefilian & Rafikov, 2019).
4 Secular behavior: simplified analytical results
We aim to elucidate the dynamical effects of the disk’s back-reaction on the planet, providing insights into the vertical structure of massive debris disks. Although Equation (15) does not admit an analytical closed-form solution for $M_{d}\neq 0$, [^5] The difficulty arises due to the non-axisymmetric component of the disk’s potential which is proportional to $I_{p}I_{d}\cos\Delta\Omega$ and may vary over time (Equation 8). In principle, an approximate analytical solution could be attempted with an educated guess for the time-averaged $I_{d}(t)$ and $\Delta\Omega(t)$, but this is not obvious to us a priori (see also Appendix B). progress can be made in a simplified case by considering only the axisymmetric component of the disk’s potential (i.e., setting $\delta B_{j,p}=0$ for all $j$), as outlined below. At the outset, we emphasize that adopting this simplified framework does not imply that the neglected terms are unimportant. Rather, it allows us to isolate and analytically examine the most basic role of disk gravity within a tractable setting, avoiding the need for numerical simulations. Nevertheless, in Appendix B, we derive an analytical expression from first principles describing the effects of the neglected terms on the planet, while their impact on the disk dynamics is discussed qualitatively in Section 6.4.
In this simplified case with $\delta B_{j,p}=0$, assuming planetesimals are initiated on coplanar orbits ($\zeta_{j}(0)=0$), Equation (15) can be solved analytically. Doing so, we find that the planetary orbit evolves such that: